# Gaia Barium Dwarfs and Their Ostensibly Ordinary Counterparts

### v1

@@ Abstract
The recently identified Gaia population of main-sequence–white dwarf (MS+WD) binaries at separations of , often with moderate eccentricities, is not readily reproduced by binary population synthesis models. Barium stars represent a closely related population whose enrichment in s-process elements both confirms the presence of a WD companion and attests to past binary interaction. It also indicates that mass transfer occurred at least during the late and post-AGB phases of the WD progenitor, when s-process elements are dredged up. In this work, we further explore the connection between the astrometrically identified Gaia MS+WD binaries and the classical barium star population. To achieve this, we used high-resolution FEROS spectroscopy to measure abundances for 30 Gaia DR3 non-single-star binaries, identifying 10 as Ba-enriched. Together with our recent analysis of archival GALAH data, this yields a sample of 38 barium dwarfs with dynamically measured WD masses, compared to only 6 previously known systems with known WD masses at these separations. We find that, in cases where metallicity is sufficiently low to facilitate efficient s-process production, barium and yttrium enrichment is often detected. This enrichment is also identified in eccentric systems, suggesting that post-AGB mass transfer mechanisms are capable of pumping eccentricity into the orbit or occur without erasing it. Our results indicate that the Gaia MS+WD binaries trace the population from which barium stars emerge. Treating the large Gaia-discovered population as an extension of known s-process enriched dwarfs opens an avenue to empirically constrain their formation and evolution.
1 Introduction
@@
2 Sample Selection and Observations
2.1 Sample Selection
@@ 3 Chemical abundances
We derived stellar parameters (, , [Fe/H]) and line-broadening velocities () from Fe lines selected for each star, following procedures adapted from S. Buder et al. (2021). Spectral fits were achieved by performing MCMC fits of pysme models to selected Fe lines, with local RV estimation and continuum detrending performed for every line segment. Seed values for , and [Fe/H] fits were taken from Gaia XP-based catalogs (R. Andrae et al., 2023; X. Zhang et al., 2023). A full description of the line selection, continuum treatment, RV corrections, and spectral fitting procedures is provided in Appendix A.
We identified 3 Ba lines (multiplets) — at 4554.03, 5853.68, 6496.90 Å in air — which are sufficiently deep and unblended with lines of other species across the range of stellar parameters of our sample. Similarly, we identified 3 Y lines — at 4900.12, 5087.42, 5200.41 Å in air. Ba and Y abundance values and their uncertainties were obtained through an MCMC analysis, with , , [Fe/H] and fixed to the previously estimated values. Local RV estimation and continuum detrending for each line segment followed the same procedure as for the Fe lines. An example of the spectral fit for Fe, Ba and Y abundances is shown in Figure 7. Due to the low SNR of the spectra, we excluded targets that did not have clearly distinguishable Ba and Y lines from the respective analysis. This excluded three targets from the Ba abundance analysis. As the Y lines are considerably shallower than the Ba lines at similar abundance levels, only 10 targets qualified for Y abundance analysis, all of them also having Ba abundance measurements.
Stellar parameters and abundances for the six two-epoch targets were derived by modeling both spectra in a joint MCMC framework, with each spectrum independently preprocessed (line selection, RV determination, continuum de-trending).
We obtained Fe and Ba abundances measured for 30 targets, 20 belonging to the (, M ) subsample, and 10 to the (M , ) subsample. Of these, 7 and 3 targets respectively have measured Y abundances. Following M. Van der Swaelmen et al. (2017) and R24, we define enrichment as cases where [Ba/Fe] or [Y/Fe] exceed 0.25 dex. A comparison of [Ba/Fe] and [Y/Fe] abundances is shown in Figure 9, confirming that Ba and Y enhancement typically occur together.
4 Results and Discussion
@@
4.1 AMRF MS+WD Systems as Ba-Stars
@@
These results indicate that the Gaia AMRF MS+WD binaries are products of interaction with AGB stars, likely through the same channels that produce Ba/CH/CEMP-s systems. Depending on metallicity, orbital period, and selection effects, they may appear as classical Ba stars. Gaia instead probes a complementary regime of shorter periods and a wider metallicity range, where enrichment is often weaker or harder to detect. Thus, even if not spectroscopically classified as s-process-enhanced, many of these systems are analogous. Notably, the current AMRF sample probes a volume more than an order of magnitude larger than the literature populations (Figure 5), providing a far broader view of the underlying parameter space. Future Gaia releases, extending to longer periods, will establish the AMRF sample as a benchmark for modeling binary interaction in intermediate-mass systems.
4.2 Eccentricity Dependence of Mass Transfer
@@
We also identify Ba enhancement in four near-circular systems hosting relatively massive WDs (M ; Figure 2b). Three of these systems have [Fe/H] dex, a regime not probed in the R24 study. We suggest that the low metallicities of these systems mitigated the effect of their progenitors’ relatively high masses, allowing sufficient Ba production for their companions to accrete and exhibit enrichment. Given the straightforward explanation for the confirmed s-process enhanced high-M systems, it is difficult to assess whether their (current) low eccentricity corresponds to higher mass transfer efficiency. Nevertheless, our sample extends the empirical evidence for mass transfer in these systems to encompass higher WD masses.
4.3 Formation Channels
@@
Our results add to a growing body of work documenting significant eccentricities in various classes of post-interaction binaries (e.g., A. I. Karakas et al., 2000; A. Jorissen et al., 2016, 2019; J. Vos et al., 2017; M. Van der Swaelmen et al., 2017; G.-M. Oomen et al., 2018; J. Müller-Horn et al., 2025). A particularly relevant parallel is provided by G.-M. Oomen et al. (2018), who studied a sample of 33 post-AGB+MS binaries with orbital periods of days, most of which exhibit significant eccentricity (). Our work demonstrates that these eccentricities persist through the post-AGB phase to the WD formation, diminishing the possibility that these systems could yet circularize via accretion and/or interaction with the post-AGB planetary nebula or circumbinary (CB) disk. Tidal evolution theory predicts that binaries in which one component fills its Roche lobe should circularize on relatively short timescales, far shorter than the expected duration of the mass transfer phase in the progenitor system (see, e.g., S. F. Portegies Zwart & F. Verbunt, 1996; G. Nelemans et al., 2001; J. F. Sepinsky et al., 2007, 2009; S. Toonen et al., 2012; S. Toonen & G. Nelemans, 2013). Multiple models have been proposed to explain this seemingly anomalous population of eccentric post-interaction systems. Broadly, these fall into two categories: (1) Mechanisms that avoid strong tidal circularization altogether; (2) Mechanisms that actively pump eccentricity, counteracting canonical tidal damping.
@@
In the first category, one possibility is that tidal forces are much weaker than expected in certain systems (e.g., J. D. Nie et al., 2017), allowing the Roche-lobe overflow (RLOF) phase to conclude before circularization can occur. This seems unlikely for our sample, except perhaps in extremely eccentric cases, because in most systems at these separations, the WD progenitor would have substantially filled its Roche lobe for much of the RGB and the entirety of the AGB phase, thereby experiencing significant tidal forces (J. P. Zahn, 1977). Another possibility is that mass transfer proceeds primarily via stellar winds, rather than RLOF, which can produce non-trivial eccentricity evolution (H. M. J. Boffin & A. Jorissen, 1988; H. M. J. Boffin, 2015; M. I. Saladino et al., 2018; M. I. Saladino & O. R. Pols, 2019; S. L. Schrøder et al., 2021). A third option is that the observed eccentricity reflects the presence of a tertiary companion, either through direct orbital perturbations or because the tertiary itself was the AGB donor, later being ejected or migrated to a wider orbit after the interaction (H. B. Perets & K. M. Kratter, 2012).
@@ 5 Outlook
We propose a survey of MS+WD systems from the catalog of S. Shahaf et al. (2024) using the HighSpec spectrograph (Y. Sofer Rimalt et al., 2024) on the upcoming Multi-Aperture Spectroscopic Telescope (MAST; Ben-Ami et al., in prep) array at Neot Smadar, Israel, with a particular focus on systems exhibiting sub-solar metallicities and moderate eccentricities. HighSpec employs a custom narrow bandpass ion-etched grating at and achieves an end-to-end efficiency of 55%, making it well suited for observing prominent spectral lines even in relatively faint targets. Our primary objective is to measure the strong Ba II absorption line at 4555 Å. The surrounding Å band also includes numerous relatively strong Fe lines, enabling precise determination of [Fe/H]. Figure 6 shows the limiting magnitude achievable at various SNRs as a function of exposure time, overlaid on the JKC -band cumulative magnitude distribution of MS+WD systems with sub-solar metallicities (as per R. Andrae et al. 2023) accessible from Israel. With up to 20 MAST telescopes operating simultaneously with HighSpec, we can observe of the brightest targets (including roughly half with ) in nights.
Appendix A Determination of RV, , , [Fe/H] and
@@
The determination of effective temperature (, surface gravity (), iron abundance ([Fe/H]) and rotational broadening velocity was carried out using a set of Fe lines custom-selected for each star. As the low SNR of our spectra makes simultaneous fitting of projected rotational velocity () and macro-turbulence velocity () challenging, we followed S. Buder et al. (2021) and fit for (defined as with set to 0) instead. The microturbulence velocity () was estimated using the empirical relation provided in the GALAH DR3 study (eqns 4 and 5; S. Buder et al., 2021). During the following procedure, where required, initial/seed values of the effective temperature (), surface gravity () and iron abundance ([Fe/H]; here used synonymously with metallicity) for each star were derived from Gaia’s low-resolution XP spectra using the XGBoost algorithm (R. Andrae et al., 2023).
@@
For each selected line, we extracted 1 Å-wide spectral segments centered on the line, merging overlapping segments. Lines were chosen based on the following criteria: (1) the SNR in the line region exceeded 10, except for three stars where the threshold was lowered to 5 due to fewer than five lines meeting the SNR condition; (2) the Fe line depth would exceed 0.4 at [Fe/H] = 0 for the star’s and ; (3) there were no lines of other species within Å exceeding 20% of the Fe line’s depth; and (4) the mean value of the continuum trend (see next) across the Å segment was greater than 0.9, with a maximum variation of less than 0.02. The fourth condition excluded regions where the blaze function caused the pipeline’s continuum normalization quality, as well as the local SNR, to drop considerably. To reduce computational expenses, when the number of such segments exceeded 30 (for 5 spectra), we selected the 30 segments with the deepest lines.
@@
To correct for the local continuum trend, we fit a smoothing spline to continuum points in 60 Å windows centered on each segment of interest using the scipy function scipy.interpolate.make_smoothing_spline. For this procedure, synthetic spectra were created with stellar parameters as described above and Doppler shifted to stellar rest frames using the nominal radial velocity (RV) determined using the H line. Continuum points were denoted as those where the model flux exceeded 0.98, which ensured a sufficient number of points for robust spline fits with negligible bias from absorption features.
@@
To determine the local RV value around each line segment, we used a 20 Å-wide region centered on the corresponding 1 Å segment. Regions (and corresponding 1 Å segments) with more than 25% of the flux points masked were discarded to ensure reliability both in RV estimation and continuum de-trending. Synthetic spectra were generated using the given stellar parameters, including only lines with absorption depths greater than 0.1. The RV was determined using cross-correlation with a built-in method from pymse. A clipping procedure was then used to identify and discard any windows with unreliable RVs, typically caused by an insufficient number of deep lines in the region. Note that these estimated RV values are not corrected for Earth’s orbital motion, as this is not required for our science case. Across all spectra, the mean (max) range in the measured RVs across all usable windows is km s-1, and the mean (max) standard deviation is km s-1.
@@
Finally, we conducted a Markov Chain Monte Carlo (MCMC) analysis to fit for , , [Fe/H] and , treating all segments simultaneously. For this we created a bank of synthetic spectra using pysme on a grid of , , [Fe/H] values, at steps of 100 K, 0.1 dex and 0.05 dex respectively, with a wavelength resolution of 0.003 Å. The line list passed to pysme during the synthesis of this bank was restricted to only Fe lines in the relevant segments in order to avoid prohibitively large computational times, as computation time scaled linearly with the number of lines in the list. This restriction does not affect our results, as the Fe lines were previously selected to avoid significant blending with lines from other species. For each star, we adopted weakly informative normal priors for and , centered on their seed values with conservative standard deviations of 250 K and 0.25 dex respectively, and uniform priors on [Fe/H] and spanning dex and km s-1 respectively. During each MCMC iteration, unbroadened synthetic spectra at the given three stellar parameters were interpolated from the grid and were rotationally and instrumentally broadened using the routines fastRotBroad, with limb-darkening parameter set at 0.81, and instrBroadGaussFast from pyastronomy (version 0.22.0; S. Czesla, 2025). Instrumental resolution at the center of each segment was computed from the wavelength solution, assuming FEROS’s sampling of 2.2 pixels per resolution element (A. Kaufer & L. Pasquini, 1998). We compared spectra broadened with pysme and pyastronomy and verified that the maximum deviation between the two was less than 1%. The broadened spectra were then interpolated onto the FEROS wavelength grid and multiplied by the continuum trend. We used as the log-likelihood function, calculated from all data points from each Fe segment selected for that star. An example of the spectral fit in Figure 7a.

### v2

@@ Abstract
The recently identified Gaia population of main-sequence–white dwarf (MS+WD) binaries at separations of , often with moderate eccentricities, is not readily reproduced by binary population synthesis models. Barium stars represent a closely related population whose enrichment in s-process elements both confirms the presence of a WD companion and attests to past binary interaction. It also indicates that mass transfer occurred at least during the late and post-AGB phases of the WD progenitor, when s-process elements are dredged up. In this work, we further explore the connection between the astrometrically identified Gaia MS+WD binaries and the classical barium star population. To achieve this, we used high-resolution FEROS spectroscopy to measure abundances for 30 Gaia DR3 non-single-star binaries, identifying 14 as Ba-enriched. Together with our recent analysis of archival GALAH data, this yields a sample of 40 barium dwarfs with dynamically measured WD masses, compared to only 6 previously known systems with known WD masses at these separations. We find that, in cases where metallicity is sufficiently low to facilitate efficient s-process production, barium and yttrium enrichment is often detected. This enrichment is also identified in eccentric systems, suggesting that post-AGB mass transfer mechanisms are capable of pumping eccentricity into the orbit or occur without erasing it. Our results indicate that the Gaia MS+WD binaries trace the population from which barium stars emerge. Treating the large Gaia-discovered population as an extension of known s-process enriched dwarfs opens an avenue to empirically constrain their formation and evolution.
I Introduction
@@
II Sample Selection and Observations
II.1 Sample Selection
@@ III Chemical abundances
We derived stellar parameters (, [Fe/H]) and line-broadening velocities (, ) from carefully selected Fe lines. Surface gravities () were fixed to values from Gaia XP–based catalogs (R. Andrae et al., 2023; X. Zhang et al., 2023), as the chosen lines are minimally pressure-broadened and therefore insensitive to . For each star, spectral fits were obtained by performing MCMC fitting of pysme models to these Fe lines, with local RV estimation and continuum detrending performed for every line segment. A full description of the Fe-line selection, continuum treatment, RV corrections, and spectral-fitting procedures is provided in Appendix A.
Ba and Y abundances were then measured using an analogous MCMC procedure, fixing the stellar parameters to the previously determined values and applying the same RV and continuum treatment. We adopted three Ba lines (4554.03, 5853.68, 6496.90 Å in air) and three Y lines (4900.12, 5087.42, 5200.41 Å in air), chosen for their depth and lack of blending across the relevant stellar parameter space. For the six two-epoch targets, stellar parameters and abundances were inferred in a joint MCMC framework, with each spectrum independently preprocessed.
Because of the low SNR of our spectra, we excluded targets without clearly distinguishable Ba and Y lines from the corresponding analysis. We obtained Fe and Ba abundances for 30 stars: 20 in the (, M ) subsample and 10 in the (M , ) subsample. Of these, Y abundances could be measured for only 7 and 3 stars, respectively, as the Y lines are considerably shallower than the Ba lines at comparable abundance levels. All fitted parameters and their uncertainties are listed in Table 1. Following M. Van der Swaelmen et al. (2017) and R24, we define enrichment as [Ba/Fe] or [Y/Fe] exceeding 0.25 dex. A comparison of [Ba/Fe] and [Y/Fe] abundances is shown in Figure 9, confirming that Ba and Y enhancement typically occur together.
IV Results and Discussion
@@
IV.1 AMRF MS+WD Systems as Ba-Stars
@@
These results indicate that the Gaia AMRF MS+WD binaries are products of interaction with AGB stars, likely through the same channels that produce Ba/CH/CEMP-s systems. Depending on metallicity, orbital period, and selection effects, they may manifest as classical Ba stars. Gaia instead probes a complementary regime of shorter periods and a wider metallicity range, where enrichment is often weaker or harder to detect. Thus, even if not spectroscopically classified as s-process-enhanced, these systems remain analogous in their overall characteristics. Notably, the current AMRF sample probes a volume an order of magnitude larger than the literature populations (Figure 5), providing a far broader view of the underlying parameter space. Future Gaia releases, extending to longer periods, will establish the AMRF sample as a benchmark for modeling binary interaction in intermediate-mass systems.
IV.2 Eccentricity Dependence of Mass Transfer
@@
We also identify high Ba enhancement in six near-circular systems hosting massive WDs (M ; Figure 2b), demonstrating that enrichment can occur even where s-process yields are expected to be suppressed due to their relatively high progenitor masses. The presence of Ba enhancement in these systems likely reflects an interplay between three factors: suppression of s-process yields with high progenitor masses, increased yields at low metallicities, and potentially enhanced accretion efficiency associated with low eccentricity. Although the current data are too sparse to disentangle these effects, these systems extend the empirical evidence for mass transfer to encompass higher-mass WDs.
IV.3 Formation Channels
@@
Our results add to a growing body of work documenting significant eccentricities in various classes of post-interaction binaries (e.g., A. I. Karakas et al., 2000; A. Jorissen et al., 2016, 2019; J. Vos et al., 2017; M. Van der Swaelmen et al., 2017; G.-M. Oomen et al., 2018; J. Müller-Horn et al., 2025). A particularly relevant parallel is provided by G.-M. Oomen et al. (2018), who studied a sample of 33 post-AGB+MS binaries with orbital periods of days, most of which exhibit significant eccentricity (). Our work demonstrates that these eccentricities persist through the post-AGB phase to the WD formation, diminishing the possibility that the Oomen at al. systems could still circularize via accretion and/or interaction with the post-AGB planetary nebula or circumbinary (CB) disk. Tidal evolution theory predicts that binaries in which one component fills its Roche lobe should circularize on relatively short timescales, far shorter than the expected duration of the mass transfer phase in the progenitor system (see, e.g., S. F. Portegies Zwart & F. Verbunt, 1996; G. Nelemans et al., 2001; J. F. Sepinsky et al., 2007, 2009; S. Toonen et al., 2012; S. Toonen & G. Nelemans, 2013). Multiple models have been proposed to explain this seemingly anomalous population of eccentric post-interaction systems. Broadly, these fall into two categories: (1) Mechanisms that avoid strong tidal circularization altogether; (2) Mechanisms that actively pump eccentricity, counteracting canonical tidal damping.
@@
In the first category, one possibility is that tidal forces are much weaker than expected in certain systems (e.g., J. D. Nie et al., 2017), allowing the Roche-lobe overflow (RLOF) phase to conclude before circularization can occur. This seems unlikely for our sample, except perhaps in extremely eccentric cases, because in most systems at these separations, the WD progenitor would have substantially filled its Roche lobe for much of the RGB and the entirety of the AGB phase, thereby experiencing significant tidal forces (J. P. Zahn, 1977). Another possibility is that mass transfer proceeds primarily via stellar winds, rather than RLOF, which can produce non-trivial eccentricity evolution (H. M. J. Boffin & A. Jorissen, 1988; H. M. J. Boffin, 2015; M. I. Saladino et al., 2018; M. I. Saladino & O. R. Pols, 2019; S. L. Schrøder et al., 2021), though the former objection may also apply here. A third option is that the observed eccentricity reflects the presence of a tertiary companion, either through direct orbital perturbations or because the tertiary itself was the AGB donor, later being ejected or migrated to a wider orbit after the interaction (H. B. Perets & K. M. Kratter, 2012).
@@ V Outlook
We propose a survey of MS+WD systems from the catalog of S. Shahaf et al. (2024) using the HighSpec spectrograph (Y. Sofer Rimalt et al., 2024) on the upcoming Multi-Aperture Spectroscopic Telescope (MAST; Ben-Ami et al., in prep) array at Neot Smadar, Israel, with a particular focus on systems exhibiting sub-solar metallicities and moderate eccentricities. HighSpec employs a custom narrow bandpass ion-etched grating at and achieves an end-to-end efficiency of 55%, making it well suited for observing prominent spectral lines even in relatively faint targets. Our primary objective will be to measure the strong Ba II absorption line at 4555 Å. The surrounding Å band also includes numerous relatively strong Fe lines, enabling precise determination of [Fe/H]. Figure 6 shows the limiting magnitude achievable at various SNRs as a function of exposure time, overlaid on the JKC -band cumulative magnitude distribution of MS+WD systems with sub-solar metallicities (as per R. Andrae et al. 2023) accessible from Israel. With up to 20 MAST telescopes operating simultaneously with HighSpec, we can observe of the brightest targets (including roughly half with ) at SNR > 20 in nights.
Appendix A Determination of RV, stellar parameters and line broadening velocities
@@
The determination of effective temperature (), Fe abundance ([Fe/H]; here used synonymously with metallicity), microturbulence velocity () and rotational broadening velocity () was carried out using a set of Fe lines custom-selected for each star. As the selected Fe lines are minimally pressure-broadened and thus largely insensitive to , surface gravity () was fixed to values obtained from Gaia XP–based catalogs (R. Andrae et al., 2023; X. Zhang et al., 2023), which also provided seed values for and [Fe/H]. As the low SNR of our spectra makes simultaneous fitting of projected rotational velocity () and macroturbulence velocity () challenging, we followed S. Buder et al. (2021) and fit for (defined as with set to 0) instead.
@@
For each selected Fe line, we extracted 1 Å-wide spectral segments centered on the line, merging overlapping segments. Lines were chosen based on the following criteria: (1) the SNR in the line region exceeded 10, except for three stars where the threshold was lowered to 5 due to fewer than five lines meeting the SNR condition; (2) the Fe line depth would exceed 0.4 at [Fe/H] = 0 for the star’s and ; (3) there were no lines of other species within Å exceeding 20% of the Fe line’s depth; and (4) the mean value of the continuum trend (see next) across the Å segment was greater than 0.9, with a maximum variation of less than 0.02. The fourth condition excluded regions where the blaze function caused the pipeline’s continuum normalization quality, as well as the local SNR, to drop considerably. To reduce computational expenses, when the number of such segments exceeded 30 (for 5 spectra), we selected the 30 segments with the deepest lines.
@@
To correct for the local continuum trend, we fit a smoothing spline to continuum points in 60 Å windows centered on each segment of interest using the scipy function scipy.interpolate.make_smoothing_spline. For this procedure, synthetic spectra were created with stellar parameters as described above and Doppler shifted to stellar rest frames using a nominal radial velocity (RV) determined using the H line. Continuum points were denoted as those where the model flux exceeded 0.98, which ensured a sufficient number of points for robust spline fits with negligible bias from absorption features.
@@
To determine the local RV value around each line segment, we used a 20 Å-wide region centered on the corresponding 1 Å segment. Regions (and corresponding 1 Å segments) with more than 25% of the flux points masked were discarded to ensure reliability both in RV estimation and continuum de-trending. Synthetic spectra were generated using the given stellar parameters, including only lines with absorption depths greater than 0.1. The RV was determined using cross-correlation with a built-in method from pymse. A clipping procedure was then used to identify and discard any windows with unreliable RVs, typically caused by an insufficient number of deep lines in the region. Across all spectra, the mean (max) range in the measured RVs across all usable windows is km s-1, and the mean (max) standard deviation is km s-1.
@@
Finally, we conducted a Markov Chain Monte Carlo (MCMC) analysis to fit for , [Fe/H], and , treating all segments simultaneously. For each star (and its corresponding fixed value), we first created a bank of synthetic spectra using pysme on a grid of , [Fe/H], and values at steps of 100,K, 0.05,dex, and 0.1,km,s-1, respectively, and with a wavelength resolution of 0.003,Å. The line list passed to pysme during this synthesis was restricted to Fe lines in the selected segments to avoid prohibitively long computation times, as runtime scaled linearly with the number of lines; this restriction does not affect our results, since the Fe lines were previously chosen to avoid significant blending with lines from other species. During each MCMC iteration, unbroadened synthetic spectra were interpolated from the precomputed grid and rotationally and instrumentally broadened using fastRotBroad (with a limb-darkening parameter of 0.81) and instrBroadGaussFast from pyastronomy (version 0.22.0; S. Czesla, 2025). The instrumental resolution at the center of each segment was determined from the wavelength solution, assuming FEROS’s sampling of 2.2 pixels per resolution element (A. Kaufer & L. Pasquini, 1998). The resulting broadened spectra were then interpolated onto the FEROS wavelength grid, multiplied by the continuum trend, and compared to the observed spectra using the standard log-likelihood function () computed from all data points across the selected Fe segments. Comparison of spectra broadened with pysme and pyastronomy confirmed that their maximum deviation was always below 1%. The MCMC posteriors were unimodal but occasionally asymmetric, especially for , and therefore we report the maximum-likelihood estimate (MLE) for each parameter, with the 68th percentile interval around the MLE serving as the uncertainty estimate. Lastly, we tested the insensitivity of the fit to by refitting [Fe/H] as a function of with the other parameters held fixed to their best-fit values, confirming that [Fe/H] remains stable within the estimated uncertainties.
Ba and Y abundances were measured using an analogous MCMC procedure, fixing the stellar parameters to the previously determined values. An example of the spectral fit to Fe, Ba and Y lines is shown in Figure 7.